# Tiling by Near Coincidence

### v1

@@ Abstract
Moiré patterns of twisted and scaled bilayers have recently emerged as a fertile source of quasiperiodic order in two-dimensional materials. Inspired by these systems, we introduce the near-coincidence method for generating quasiperiodic tilings of the plane. The method is intuitive—admitting pairs of nearly coincident points from superimposed layers—yet rigorous, as it maps naturally to the well-established cut-and-project formalism. It reproduces classical tilings, including the Ammann–Beenker, the Niizeki–Gähler, and the square and hexagonal Fibonacci tilings. It also uncovers new tilings not likely to arise in conventional constructions, with relative frequencies of local configurations that may take transcendental values. The near-coincidence method is algorithmically simple and already realized in an application that generates tilings from specified layer parameters and coincidence conditions. Future extensions include trilayer systems, where preliminary results yield dodecagonal order with square layers, and very small twist angles, where the method may capture the giant moiré patterns of bilayer and trilayer graphene.
@@ I Statement of the problem and physical motivation
Quasiperiodic tilings of the plane—like the famous octagonal (8-fold) Ammann–Beenker tiling [Beenker82, Ammann92], decagonal (10-fold) Penrose tiling [Penrose74], and dodecagonal (12-fold) Stampfli [Stampfli86] and Niizeki–Gähler [Niizeki87, Gahler88] tilings—have, for decades, served as canonical mathematical models for two-dimensional quasicrystals, and for the high-symmetry surfaces of three-dimensional quasicrystals. The challenge of constructing such quasiperiodic tilings, with prescribed rotational symmetry, lies in the physical requirement that the collection of vertices form a Delone set—that is—be both uniformly discrete (so that vertices are never arbitrarily close to one another) and relatively dense (so that no arbitrarily large voids appear).
A naive construction, taking the closure under vector addition of an $n$-fold symmetric star of vectors, yields a valid two-dimensional tiling only when $n=3,4,$ or $6$, in which case the outcome is periodic. For all other values of $n>2$ the point set becomes dense in the plane, thereby violating uniform discreteness. Thus, to achieve a valid Delone set without compromising the rotational symmetry, one must employ a systematic selection procedure that decides which of the candidate points are retained and which must be excluded.
Two rigorous and well-established procedures for doing precisely this are the dual-grid and the cut-and-project methods, both thoroughly presented in standard textbooks [Senechal95, Baake13]. In the dual-grid method, the acceptance of vertices into the tiling—and of points into the associated Delone set—is governed by the topology of an appropriately defined grid in a dual space. In the cut-and-project method, by contrast, the closure of the initial $n$-fold symmetric star of vectors is embedded in a higher-dimensional space, where it forms a periodic lattice. Points of this lattice are projected back onto physical space, but only if their projection, onto the orthogonal complement subspace, falls within a well-defined acceptance domain, thus producing the quasiperiodic tiling.
In the present work we introduce a new and more intuitive procedure for generating quasiperiodic tilings of the plane, inspired by the recent surge of interest in twisted bilayer graphene [Andrei20]—which is known to exhibit long-range aperiodic order with hexagonal [Lai25] and dodecagonal [Ahn18] symmetry—as well as in other multilayered two-dimensional materials. These physical systems continue to attract attention because of their remarkable tunability, which allows one to engineer properties such as superconductivity [Bistritzer11, Cao18], controllable magnetism [Gonzalez17], and ferroelectricity [Vizner21], together with correlated electronic behavior that clearly reflects the underlying quasiperiodic order [Uri23].
@@
When the rotation angle is commensurate, or alternatively, the scaling factor is rational, the bilayer admits an infinite set of perfect coincidence sites. These sites form a periodic sublattice and hence a periodic tiling with an enlarged unit cell, preserving the symmetry of the individual layers up to a reorientation of the point group. A thorough analysis of such commensurate bilayers has recently been provided by Gratias and Quiquandon [Gratias23]. However, in the generic, incommensurate case—analyzed more recently by the same authors [Quiquandon25]—at most a single perfect coincidence site can exist. Nevertheless, for any threshold $r>0$, there is always a countable infinity of pairs of nearly coincident points—one from each layer—that are separated by no more than a distance $r$. We argue here that these points of near coincidence provide a natural and intuitive basis for constructing a quasiperiodic tiling with the required symmetry.
To emphasize the intuitiveness of this near-coincidence method, we begin in Sec. II with a step-by-step illustration of the construction of an octagonal tiling from two superimposed square lattices rotated by $45^{\circ}$ (cf. Fig. 1a). In Sec. III we establish the validity of the method by mapping it rigorously to the cut-and-project formalism, and by revealing the substitution rules—originally derived by Beenker [Beenker82]—by merely varying the coincidence threshold. We then demonstrate the versatility of the approach in Sec. IV, generating dodecagonal tilings [Niizeki87, Gahler88] from bilayers of triangular lattices rotated by $30^{\circ}$ (cf. Fig. 1b), and in Sec. V, producing quasiperiodic square [Lifshitz02] and hexagonal [Coates24] Fibonacci tilings from bilayers scaled by the golden mean (cf. Fig. 1c). We conclude in Sec. VI with a summary and an outlook toward future directions.
@@ II Generating a Tiling Using the Near-Coincidence Method
The guiding principle is straightforward. Pairs of points that lie closer together are more likely to contribute a vertex to the tiling, whereas more widely separated pairs are less likely to do so. The task, therefore, is (A) to define a systematic criterion for deciding which near-coincidence sites are accepted, mapping them consistently into a point set of potential tiling-vertices that satisfy the Delone conditions; (B) to connect these points with appropriately chosen edges; and finally (C) to perform local cleaning, removing excess points or crossed edges according to simple local rules, in order to obtain a proper tiling.
@@ II.2 Connecting Vertices with Edges
Once the potential tiling-vertices have been selected, one must decide how to connect them with edges. There are several seemingly arbitrary choices at this stage, which may be influenced by physical or aesthetic considerations. First, the overall scale of edge lengths is imposed by the choice of threshold $r$, which controls the vertex density. Thus, a smaller $r$ yields longer edges, while a larger $r$ yields shorter edges. In addition, special choices of $r$ may allow a minimal number of permitted edge lengths—possibly only one, as in the octagonal example of Fig. 2—while other choices may require additional edge lengths.
Once the allowed edge lengths are determined, one begins at the origin and extends outward, drawing edges between all pairs of vertices separated by the permitted edge lengths. This procedure, applied to the octagonal point set using a single allowed edge length, yields only two prototiles—a square and a $45^{\circ}$-rhomb—as shown in Fig. 2a.
@@ II.3 Cleaning Up the Tiling
Figure 4: Octagonal tilings obtained by the near-coincidence method as described in Sec. II. (a) The octagonal tiling obtained by admitting all the vertices, shown in Fig. 2b, whose coincidence falls within the brown circular coincidence window, depicted in panel (c). The tiling requires three different edge lengths, and contains four prototiles—a square, a $45^{\circ}$-rhomb, a trapezoid, and a kite—displayed beneath the tiling. (b) The octagonal tiling obtained when insisting on a single edge length, thus removing the weaker-coincidence vertex in every pair of nearby vertices. This removes excess points, whose positions are indicated by empty circles, and leads to a tiling containing only the square and the $45^{\circ}$-rhomb prototiles, displayed beneath the tiling. As demonstrated in Sec. III, the vertices that are accepted into this tiling correspond to using the light-brown octagonal coincidence window, shown in panel (c), inscribed within the original circular one. This is the well-known Ammann–Beenker [Beenker82, Ammann92] tiling. (c) Coincidence windows selecting the point sets that are admitted into the octagonal tilings in (a) and (b).
@@
If there is insufficient space for two atoms at such nearby positions, one must eliminate one of the vertices, splitting the encompassing stretched hexagon into a single square and a pair of rhombs, the two options of which are shown in Figs. 3c and 3d. The natural approach is to prefer the vertex that originates from higher coincidence, corresponding to the closer red–blue pair, and to discard the lower-coincidence vertex. This choice is shown in Fig. 3d. A patch of the final octagonal tiling, after removing all excess points, is shown in Fig. 4b. The informed reader may recognize that this tiling looks very much like the well-known Ammann–Beenker tiling [Beenker82, Ammann92]. We establish in the following section that this is indeed the case. It should be noted that redundancies in local tile configurations, associated with small variations in vertex positions like the ones in Figs. 3c and 3d, are not mere mathematical artifacts, but are encountered regularly in physical quasicrystals, where they correspond to so-called phason flips [Lifshitz11]. These are local rearrangements in which atoms may fluctuate between nearly degenerate sites due to thermal excitation, as observed in high-resolution transmission electron microscopy of metallic-alloy quasicrystals [Edagawa00].
@@ III.1 Mapping to The Cut-and-Project Method
Figure 5 displays such a mapping of about a thousand tiling vertices. Each vertex is mapped from its position $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{p}_{1}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{p}_{2}})/2$ in the tiling to the position of its corresponding separation vector ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{p}_{1}}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{p}_{2}}$, relative to the center of a circular coincidence window. The color coding is maintained for the accepted vertices, while the discarded excess vertices are drawn in black. The distribution of the excluded points outlines, with striking clarity, a perfect octagonal coincidence window.
This rather profound observation implies that from the outset, one could have chosen a polygonal coincidence window, instead of a trivial isotropic one, thereby excluding all the excess points in advance, and avoiding the local-rule clean-up step. Thus, what appeared in Sec. II as an ad hoc pragmatic act of local cleaning can be replaced by an exact global criterion. Of course, one needs to be careful in choosing a proper polygon that does not break the original symmetry of the bilayer. If one uses a square coincidence window, for example, one obtains a tetragonal, or square, quasiperiodic tiling, instead of the octagonal one.
@@
At first sight this may seem like a convenience. Yet, up to a factor of 2, it coincides exactly with one of the standard so-called star maps, used in the cut-and-project formalism to decide whether a particular integer linear combination of an 8-fold symmetric star of vectors belongs in the tiling (see, for example, Fig. 3a of Socolar [Socolar89]). The star map associates with every potential vertex $\mathbf{p}$}, of the Ammann–Beenker tiling, a unique two-dimensional orthogonal complement, $\mathbf{p}$$^{\star}$, given by
@@
The vertex $\mathbf{p}$} is accepted into the tiling if its star map $\mathbf{p}$$^{\star}$ belongs to the acceptance domain AD. Thus, the tiling defined in Eq. (1) is equivalent, to within a factor of 2, to the tiling defined using the star map,
$\mathcal{T}=\left\{\tfrac{1}{2}\sum_{k=0}^{3}n_{k}\mathbf{a}_{k}\;\middle|\;\sum_{k=0}^{3}n_{k}\mathbf{b}_{k}\in\text{AD}\right\}.$ (7)
The star map of the cut-and-project method finds a natural reinterpretation in the near-coincidence framework as a direct distance measure between paired points in the bilayer. The coincidence window is nothing other than a real-space manifestation of the acceptance domain. This equivalence—first hinted at by the outline of the discarded points—confirms that our intuitive near-coincidence method reproduces, in full rigor, the well-known Ammann–Beenker construction. As an additional observation, we have discovered that replacing the octagonal acceptance domain of the Ammann–Beenker tiling by its inscribing circle yields a new octagonal tiling with square, $45^{\circ}$-rhomb, kite, and trapezoid prototiles.
@@ III.2 Substitution Rules
In this way, the known substitution rules of the Ammann–Beenker tiling [Beenker82, Figs. 6.4 and 6.5] emerge naturally from the near-coincidence framework. Figure 7b shows them schematically: the rhomb subdivides symmetrically into three smaller rhombs and four half-squares, whereas the square subdivides asymmetrically into a central square surrounded by four rhombs and four half-squares. The alternating orientations of the squares, indicated by red arrows, can be read directly from the geometry of successive inflated generations in Fig. 7a. This provides yet another confirmation that the near-coincidence method faithfully reproduces the well-known standard constructions.

### v2

@@ Abstract
Moiré patterns of twisted and scaled bilayers have recently emerged as a fertile source of quasiperiodic order in two-dimensional materials. Inspired by these systems, we introduce the near-coincidence method for generating quasiperiodic tilings of the plane. The method is intuitive—admitting pairs of nearly coincident points from superimposed layers—yet rigorous, as it maps naturally to the well-established cut-and-project formalism. It reproduces classical tilings, including the Ammann–Beenker, the Niizeki–Gähler, and the square and hexagonal Fibonacci tilings, and also reveals new tilings that are unlikely to arise from conventional constructions. The near-coincidence method is algorithmically simple and already realized in a web-based application that generates tilings from specified layer parameters and coincidence conditions. Future extensions include trilayer systems, where preliminary results yield dodecagonal order with square layers, and very small twist angles, where the method may capture the giant moiré patterns of bilayer and trilayer graphene.
@@ I Statement of the problem and physical motivation
Quasiperiodic tilings of the plane—like the famous octagonal (8-fold) Ammann–Beenker tiling [1, 2], decagonal (10-fold) Penrose tiling [3], and dodecagonal (12-fold) Stampfli [4] and Niizeki–Gähler [5, 6] tilings—have, for decades, served as canonical mathematical models for two-dimensional quasicrystals, and for the high-symmetry surfaces of three-dimensional quasicrystals. The challenge of constructing such quasiperiodic tilings, with prescribed rotational symmetry, lies in the physical requirement that the collection of vertices form a Delone set—that is—be both uniformly discrete (so that vertices are never arbitrarily close to one another) and relatively dense (so that no arbitrarily large voids appear).
A naive construction, taking the closure under vector addition of an $n$-fold symmetric star of vectors, yields a valid two-dimensional tiling only when $n=3,4,$ or 66, in which case the outcome is periodic. For all other values of $n>2$ the point set becomes dense in the plane, thereby violating uniform discreteness. Thus, to achieve a valid Delone set without compromising the rotational symmetry, one must employ a systematic selection procedure that decides which of the candidate points are retained and which must be excluded.
Two rigorous and well-established procedures for doing precisely this are the dual-grid and the cut-and-project methods, both thoroughly presented in standard textbooks [7, 8]. In the dual-grid method, the acceptance of vertices into the tiling—and of points into the associated Delone set—is governed by the topology of an appropriately defined grid in a dual space. In the cut-and-project method, by contrast, the closure of the initial $n$-fold symmetric star of vectors is embedded in a higher-dimensional space, where it forms a periodic lattice. Points of this lattice are projected back onto physical space, but only if their projection, onto the orthogonal complement subspace, or internal space, falls within a well-defined acceptance domain, thus producing the quasiperiodic tiling.
In the present work we introduce a new and more intuitive procedure for generating quasiperiodic tilings of the plane, inspired by the recent surge of interest in twisted bilayer graphene [9]—which is known to exhibit long-range aperiodic order with hexagonal [10] and dodecagonal [11] symmetry—as well as in other multilayered two-dimensional materials. These physical systems continue to attract attention because of their remarkable tunability, which allows one to engineer properties such as superconductivity [12, 13], controllable magnetism [14], and ferroelectricity [15], together with correlated electronic behavior that clearly reflects the underlying quasiperiodic order [16].
@@
When the rotation angle is commensurate, or alternatively, the scaling factor is rational, the bilayer admits an infinite set of perfect coincidence sites. These sites form a periodic sublattice and hence a periodic tiling with an enlarged unit cell, preserving the symmetry of the individual layers up to a reorientation of the point group. A thorough analysis of such commensurate bilayers has recently been provided by Gratias and Quiquandon [17]. However, in the generic, incommensurate case—analyzed more recently by the same authors [18]—at most a single perfect coincidence site can exist. Nevertheless, for any threshold $r>0$, there is always a countable infinity of pairs of nearly coincident points—one from each layer—that are separated by no more than a distance $r$. We argue here that these points of near coincidence provide a natural and intuitive basis for constructing a quasiperiodic tiling with the required symmetry.
To emphasize the intuitiveness of this near-coincidence method, we begin in Sec. II with a step-by-step illustration of the construction of an octagonal tiling from two superimposed square lattices rotated by $45^{\circ}$ (cf. Fig. 1a). Informed readers may skip this part and proceed directly to Sec. III, where we establish the validity of the method by mapping it to the cut-and-project formalism, and by revealing the substitution rules—originally derived by Beenker [1]—by varying the coincidence threshold. We then demonstrate the versatility of the approach in Sec. IV, generating dodecagonal tilings [5, 6] from bilayers of triangular lattices rotated by $30^{\circ}$ (cf. Fig. 1b), and in Sec. V, producing quasiperiodic square [19] and hexagonal [20] Fibonacci tilings from bilayers scaled by the golden mean (cf. Fig. 1c). We conclude in Sec. VI with a summary and an outlook toward future directions.
@@ II Generating a Tiling Using the Near-Coincidence Method
The guiding principle is straightforward. Pairs of points that lie closer together are more likely to contribute a vertex to the tiling, whereas more widely separated pairs are less likely to do so. The task, therefore, is (A) to define a systematic criterion for deciding which near-coincidence sites are accepted, mapping them consistently into a point set of potential tiling-vertices that satisfy the Delone conditions; (B) to connect these points with appropriately chosen edges; and finally (C) to perform local cleaning, removing excess points or crossed edges according to simple local rules, in order to obtain a proper tiling  [^1] We note that the latter two steps appear in other tiling constructions, most commonly when one starts with a smooth quasiperiodic field and then identifies a discrete tiling by (A) selecting maxima as potential vertices; (B) connecting them with edges; and (C) resolving local ambiguities. See, for example, [36, 37, *Mace16, 39]..
@@ II.2 Connecting Vertices with Edges
Once the potential tiling-vertices have been selected, one must decide how to connect them with edges. There are several seemingly arbitrary choices at this stage, which may be influenced by physical or aesthetic considerations. For example, when modeling a physical system within a tight-binding framework, the choice of edge connections may become important, as it can affect both the topology and the relative strengths of hopping processes between nearby sites.
First, the overall scale of edge lengths is imposed by the choice of threshold $r$, which controls the vertex density. Thus, a smaller $r$ yields longer edges, while a larger $r$ yields shorter edges. In addition, special choices of $r$ may allow a minimal number of permitted edge lengths—possibly only one, as in the octagonal example of Fig. 2—while other choices may require additional edge lengths.
Once the allowed edge lengths are determined, our default, empirically motivated approach is to draw edges between all pairs of vertices separated by the permitted edge lengths. Applied to the octagonal point set with a single allowed edge length, this procedure produces only two prototiles—a square and a $45^{\circ}$-rhomb—as shown in Fig. 2a. Of course, one may also obtain a tiling from the selected point set by a Voronoi diagram, or by its dual Delone triangulation [8, sec. 2.2], but these generic constructions do not in general recover the particular edge set or prototile structure of interest here.
@@ II.3 Cleaning Up the Tiling
Figure 4: Octagonal tilings obtained by the near-coincidence method as described in Sec. II. (a) The octagonal tiling obtained by admitting all the vertices, shown in Fig. 2b, whose coincidence falls within the brown circular coincidence window, depicted in panel (c). The tiling requires three different edge lengths, and contains four prototiles—a square, a $45^{\circ}$-rhomb, a trapezoid, and a kite—displayed beneath the tiling. (b) The octagonal tiling obtained when insisting on a single edge length, thus removing the weaker-coincidence vertex in every pair of nearby vertices. This removes excess points, whose positions are indicated by empty circles, and leads to a tiling containing only the square and the $45^{\circ}$-rhomb prototiles, displayed beneath the tiling. As demonstrated in Sec. III, the vertices that are accepted into this tiling correspond to using the light-brown octagonal coincidence window, shown in panel (c), inscribed within the original circular one. This is the well-known Ammann–Beenker [1, 2] tiling. (c) Coincidence windows selecting the point sets that are admitted into the octagonal tilings in (a) and (b).
@@
If there is insufficient space for two atoms at such nearby positions, one must eliminate one of the vertices, splitting the encompassing stretched hexagon into a single square and a pair of rhombs, the two options of which are shown in Figs. 3c and 3d. The natural approach is to prefer the vertex that originates from higher coincidence, corresponding to the closer red–blue pair, and to discard the lower-coincidence vertex. This choice is shown in Fig. 3d. A patch of the final octagonal tiling, after removing all excess points, is shown in Fig. 4b. The informed reader may recognize that this tiling looks very much like the well-known Ammann–Beenker tiling [1, 2]. We establish in the following section that this is indeed the case.
It should be noted that redundancies in local tile configurations, associated with small variations in vertex positions like the ones in Figs. 3c and 3d, are not mere mathematical artifacts, but are encountered regularly in physical quasicrystals, where they correspond to so-called phason flips [22]. These are local rearrangements in which atoms may fluctuate between nearly degenerate sites due to thermal excitation, as observed in high-resolution transmission electron microscopy of metallic-alloy quasicrystals [23].
@@ III.1 Mapping to The Cut-and-Project Method
The expert reader may immediately notice that the near-coincidence construction admits a natural interpretation within a generalized version of the cut-and-project procedure, mentioned briefly in Sec. I. In this view, one considers the direct product of the two lattices defining the bilayer, and regards pairs of points $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{p}_{1}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{p}_{2}})$ as elements of a higher-dimensional lattice. The physical-space coordinates are then obtained by a projection that combines the two layers—for example, by taking their midpoint—while the internal-space coordinates are given by their difference ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{p}_{1}}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{p}_{2}}$, whose restriction to a bounded acceptance domain is defined by the coincidence window. In this more general setting, the physical and internal spaces need not be orthogonal, as is often assumed in standard presentations of the cut-and-project method.
At the same time, the formulation adopted here emphasizes the physical origin of the construction, starting from real-space bilayers and their near coincidences, rather than from an abstract higher-dimensional embedding. In what follows, we show explicitly how the near-coincidence method leads directly to the standard cut-and-project description in the octagonal case.
@@
Figure 5 displays such a mapping of about a thousand tiling vertices. Each vertex is mapped from its position $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{p}_{1}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{p}_{2}})/2$ in the tiling to the position of its corresponding separation vector ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{p}_{1}}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{p}_{2}}$, relative to the center of a circular coincidence window. The color coding is maintained for the accepted vertices, while the discarded excess vertices are drawn in black. The distribution of the excluded points outlines, with notable clarity, a perfect octagonal coincidence window.
This observation suggests that from the outset, one could have chosen a polygonal coincidence window, instead of a trivial isotropic one, thereby excluding all the excess points in advance, and avoiding the local-rule clean-up step. Thus, what appeared in Sec. II as an ad hoc pragmatic act of local cleaning can be replaced by an exact global criterion. Of course, one needs to be careful in choosing a proper polygon that does not break the original symmetry of the bilayer. If one uses a square coincidence window, for example, one obtains a tetragonal, or square, quasiperiodic tiling, instead of the octagonal one.
@@
At first sight this may seem like a convenience. Yet, up to a factor of 2, it coincides exactly with one of the standard so-called star maps, used in the cut-and-project formalism to decide whether a particular integer linear combination of an 8-fold symmetric star of vectors belongs in the tiling (see, for example, Fig. 3a of Socolar [24]). The star map associates with every potential vertex $\mathbf{p}$}, of the Ammann–Beenker tiling, a unique two-dimensional orthogonal complement, $\mathbf{p}$$^{\star}$, given by
@@
[…]
The star map of the cut-and-project method finds a natural reinterpretation in the near-coincidence framework as a direct distance measure between paired points in the bilayer. The coincidence window is therefore a real-space manifestation of the acceptance domain. This equivalence—first hinted at by the outline of the discarded points—shows that the near-coincidence method reproduces, in a direct and transparent way, the well-known Ammann–Beenker construction.
As an additional observation, we have discovered that replacing the octagonal acceptance domain of the Ammann–Beenker tiling by its inscribing circle yields a new octagonal tiling with square, $45^{\circ}$-rhomb, kite, and trapezoid prototiles. We should note that circular or nonpolygonal windows have been considered in extensions of the cut-and-project method [25]; here, they arise naturally as the simplest choice within the bilayer perspective.
@@ III.2 Substitution Rules
In this way, the known substitution rules of the Ammann–Beenker tiling [1, Figs. 6.4 and 6.5] emerge naturally from the near-coincidence framework. Figure 7b shows them schematically: the rhomb subdivides symmetrically into three smaller rhombs and four half-squares, whereas the square subdivides asymmetrically into a central square surrounded by four rhombs and four half-squares. The alternating orientations of the squares, indicated by red arrows, can be read directly from the geometry of successive inflated generations in Fig. 7a. This provides yet another confirmation that the near-coincidence method faithfully reproduces the well-known standard constructions.